OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 25 degrees Celsius is.

OpenStudy (anonymous):

\[C = \frac{ 5 }{ 9 } (F - 32)\]

OpenStudy (anonymous):

multiply both sides by 9 9C = 5(F-32) divide both sides by 5 9C/5 = F-32 =>F = 9C/5 +32-----> part a put C=25 => F =9(25)/5 +32 => F =45+32 => F =77 -----> part b